OpenStudy (anonymous):

How do I solve this system of linear equation using the substitution method? 15x-y=14 3x-4y=18

OpenStudy (linknissan):

first equation... bring y to the right and 14 to the left.. its 15x-14=y

OpenStudy (linknissan):

now put the value of y in the second equation

OpenStudy (anonymous):

http://www.purplemath.com/modules/systlin4.htm

OpenStudy (anonymous):

Right, which would lead me to 3x-4(15x-14)=18 3x-60x+56=18 57x=-38

OpenStudy (linknissan):

yes....

OpenStudy (anonymous):

-57x=-38 x=2/3 subbing into (2): 3(2/3)+4y=18 2+4y=18 y=16/4=4 solution: (2/3, 4)